Last year Tammy had $4,500 to invest in two savings accounts. One account earns 7% simple annual interest, and the other account earns 6% simple annual interest. At the end of the year, Tammy had earned a total of $290 in interest from the two savings accounts.
Part A Write a system of two equations that could be used to find the amount of money invested in each account. Let x represent the amount invested at 7%, and let y represent the amount invested at 6%.
Part B Solve the system of equations. How much did Tammy have invested at each rate?

Found 2 solutions by addingup, jorel555:
Answer by addingup(3677)

0.07x+0.06(4500-x) = 290
0.07x-0.06x+270 = 290
0.01x = 20
x = 2,000
$2000 are invested at 7% and 4500-2000 = $2500 are at 6%

Answer by jorel555(1290)

x+y=4500
.07x+.06y=290
---------------
7x+6y=29000
6x+6y=27000;
x=2000 invested at 7%
y=2500 invested at 6%!!!!!!!!!!